If the total bill at a restaurant, including a 15% tip, is
$65.32, what was the cost of the meal alone?

Answer by mukhopadhyay(490)

If the cost of the meal is x dollars, x + 15% of x = 65.32
x +15% of x = (1.15)x = 65.32
=> x = 65.32/15 = 56.80;
The cost of the meal alone is $56.80